a=i+j+k ; b=2i=k; c=2i+j+k; d=i+j+2k then find [a*b].[c*d]

Grade:12

1 Answers

Arun
6 years ago
Dear student
 
First find 
A X B = (i + j+ k) X (2i -k) = (j -2k -i +2j) = (-i +3j -2k)
Now
C X D = (2i + j+ k) X (i + j +2k) = (2k -4j -k + 2i +j -i) = (i-3j +k)
Now we have to find dot product
Hence answer = -1 -9 -2 = -12
 
 
Regards
Arun (askIITians forum expert)